snajay wants to earn extra money for college. Because he like baking, he decided to sell chocolate chip cookies on campus. The cost to start his business s $120 and each cookie costs 10 cents to make. He will sell the cookies for 50 cents each. How many cookies does he need to sell to break even?
I am a little confused with this question, since we are learning to develop two equations to find the point of intersection. Some help would be great please! I cant figure out what the two equations are supposed to be.

Answer by math_helper(2461)

On the one hand there is a cost equation. This represents the money he needs to spend (in dollars) to make x cookies:
C(x) = 120 + (0.1)x

Now when he sells cookies, there is a revenue equation for x cookies sold. This represents the money he takes in:
R(x) = (0.5)x

He breaks even when R(x) = C(x):
(0.5)x = 120+(0.1)x
(0.4)x = 120
x = 120/(0.4) = 300

You can also plot R(x) and C(x) to see where they intersect.